# STREAMS: An Assistive Multimodal AI Framework for Empowering Biosignal Based Robotic Controls

### v1

@@ Abstract
End-effector based assistive robots face persistent challenges in generating smooth and robust trajectories when controlled by human’s noisy and unreliable biosignals such as muscle activities and brainwaves. The produced endpoint trajectories are often jerky and imprecise to perform complex tasks such as stable robotic grasping. We propose STREAMS (Self-Training Robotic End-to-end Adaptive Multimodal Shared autonomy) as a novel framework leveraged deep reinforcement learning to tackle this challenge in biosignal based robotic control systems. STREAMS blends environmental information and synthetic user input into a Deep Q Learning Network (DQN) pipeline for an interactive end-to-end and self-training mechanism to produce smooth trajectories for the control of end-effector based robots. The proposed framework achieved a high-performance record of 98% in simulation with dynamic target estimation and acquisition without any pre-existing datasets. As a zero-shot sim-to-real user study with five participants controlling a physical robotic arm with noisy head movements, STREAMS (as an assistive mode) demonstrated significant improvements in trajectory stabilization, user satisfaction, and task performance reported as a success rate of 83% compared to manual mode which was 44% without any task support. STREAMS seeks to improve biosignal based assistive robotic controls by offering an interactive, end-to-end solution that stabilizes end-effector trajectories, enhancing task performance and accuracy.
@@ I INTRODUCTION
However, such individualized biosignal based control approaches and their associated decoders are not able to generate stable outputs (trajectories) particularly when facing the challenge of continuous control of an end-effector based system (e.g. computer cursor) for accurate target acquisition (Fig 1)[11, 12]. This challenge could be more destructive when controlling the endpoint of a physical robotic system where a minimum jerky motion causes failure in tasks such as robotic grasping for a patient with severe motor impairment [14]. Additionally, the required high level of engagement for continuous control of traveled trajectory imposes a high cognitive load on most patients particularly those with brain implants [12]. This cognitive burden was further escalated by the uncertainty of human intention decoders in real-world scenarios of human-robot collaboration for target acquisition, as studied by Dani et al. [15]. Most of the previous research studies on shared human-robot control [16, 17, 18, 19] assumed known or explicitly predicted targets, lacking real-time responsiveness to dynamic user intentions, demands, or changes in the environment’s structure. To this end, we present STREAMS (Self-Training Robotic End-to-end Adaptive Multimodal Shared autonomy), a novel framework using deep reinforcement learning along with a DQN approach, to overcome the barriers of current human-robot shared control approaches in reach-to-grasp objects in 2D workspace by focusing on the following highlighted contributions:
@@ III-A Problem formulation and framework
In our case study, we consider $\tilde{h_{t}}\in\{-1,0,1\}$, which represents decoded values of 1D head movements corresponding to left, neutral, and right, respectively. The action space $\mathcal{A}$ consists of five possible actions: forward, backward, right, left, and hold. We focus on controlling the end-effector position directly, while the positions of other joints in the robotic arm are determined using inverse kinematics. The Q-value function $Q(s_{t},a)$ is approximated by a DQN, which integrates convolutional layers to process image inputs, embedding layers to represent human inputs, and a late fusion mechanism to combine information from both modalities. The output is passed through linear layers to make the final action decision as it illustrated in Fig 3.
@@
Here, the function $\text{Flip}(h^{ideal}_{t})$ randomly changes $h^{ideal}_{t}$ to one of the other two possible values $\{-1,0,1\}$. The noise generation process presented here for our head movement signal is oriented to the characteristics of EEG signals, which normally have a lower SNR compared with other biosignals [25]. Different studies [26, 27, 28, 29] show that for the multiclass motor imagery classification problem using EEG, typical decoder accuracies range between 70% and 90%. The same variability in decoder outputs is emulated here by considering three different levels of noise assigned probabilities ($p$) of 0.2, 0.3, and 0.4 during the training phase. The pseudo-code for the training phase of the algorithm is presented in Algorithm 1. The DQN in the STREAMS framework is trained using experience replay, where past experiences $(s_{t},a_{t},r_{t},s_{t+1})$ are stored in a replay buffer, and mini-batches are sampled to update the network parameters $\theta$. A target network with parameters $\theta^{\prime}$ is used to stabilize learning. The loss function involves the difference between the predicted and target Q-values, incorporating the reward $r$ and discount factor $\gamma$. The performance is evaluated on metrics such as task success rate, trajectory efficiency, and robustness to input noise. STREAMS is trained in a simulated environment with a hyperparameter random search using the Adam optimizer and an $\epsilon$-greedy strategy for exploration, with $\epsilon$ decaying from 0.9 to 0.1. Training includes 25,000 episodes and randomization of cup positions to prevent overfitting and enhance generalization across tasks.
@@ III-B User study design
This IRB-approved study involved five healthy participants aged 22-36 years. The experimental setup included a Kinova Jaco robotic arm, an RGB camera for workspace perception, an Inertial Measurement Unit (IMU)-based head movement detector, and two randomly selected target objects which are supposed to be grasped by user inputs (see Fig. 4). The head movement detector measured the angle of the head toward the sides and decoded these movements into three discrete control inputs: right, left, and neutral. These user inputs was used to guide the robotic arm to reach two different target cups. A customized YOLOv8 [30] was trained for real-time object detection and distance tracking. The framework was generalized to handle a wide range of target objects including different types of cups, bottles, and containers. The reach-to-grasp operation in a planar 2D space required comparing Manual (pure head movement control) and Assistive (with STREAMS) modes. Each participant completed 40 trials: 20 in Manual mode and 20 in Assistive mode, with 10 trials for each target (left and right) in both modes. target positions remained constant throughout the experiment to have a systematic user study across participants. In this case study, users controlled lateral motion through head movements, while the end-effector constantly moved 2 cm forward with each step. Only a noise level of 0.3 ($p$ in 5) was added to simulate biosignal control for our user study. Grasping was automatically triggered whenever the end-effector was in proximity to the target sufficiently.
Figure 4: Components of the STREAMS framework user study
Before data collection, participants were trained in both modes and completed practice trials. Mode order was randomized to minimize learning effects. Data collection included video recording, trajectory extraction, targets and end-effector position tracking, and success rate recording. Participants also completed questionnaires on convenience, effort, and satisfaction for each mode. Performance metrics compared trajectory qualities, success rates, and user feedback between Manual and Assistive modes. A paired t-test was used to determine statistical significance of observed differences.
@@ V Conclusion
This study introduced STREAMS as a task support framework for shared autonomy in assistive robotics, addressing the challenge of generating robust trajectories from unreliable and noisy user inputs for precise robotic grasping. Our generalizable framework incorporates several key components, including an end-to-end self-learning policy for interaction with dynamic user intent and adaptation to environmental changes. STREAMS demonstrated the potential for generating smooth and stable trajectories across various noise levels and human users, without relying on pre-existing datasets. The efficacy of the proposed framework was validated in both simulated environments and real user cases, showing improvements in task performance and accuracy compared to manual control with no task support. Future research directions could include expanding the framework’s capabilities to handle more complex environments and tasks.

### v2

@@ Abstract
End-effector based assistive robots face persistent challenges in generating smooth and robust trajectories when controlled by human’s noisy and unreliable biosignals such as muscle activities and brainwaves. The produced endpoint trajectories are often jerky and imprecise to perform complex tasks such as stable robotic grasping. We propose STREAMS (Self-Training Robotic End-to-end Adaptive Multimodal Shared autonomy) as a novel framework leveraged deep reinforcement learning to tackle this challenge in biosignal based robotic control systems. STREAMS blends environmental information and user input into a Deep Q Learning Network (DQN) pipeline for an interactive end-to-end and self-training mechanism to produce smooth trajectories for the control of end-effector based robots. The proposed framework achieved a high-performance record of 98% in simulation with dynamic target estimation and acquisition without any pre-existing datasets. As a zero-shot sim-to-real user study with five participants controlling a physical robotic arm with noisy head movements, STREAMS (as an assistive mode) demonstrated significant improvements in trajectory stabilization, user satisfaction, and task performance reported as a success rate of 83% compared to manual mode which was 44% without any task support. STREAMS seeks to improve biosignal based assistive robotic controls by offering an interactive, end-to-end solution that stabilizes end-effector trajectories, enhancing task performance and accuracy. The STREAMS codes and demo videos can be accessed at: https://github.com/AbiriLab/STREAMS
Index Terms:
Shared autonomy, biosignal-based control, multimodal interfaces, self-learning algorithms.
@@ I INTRODUCTION
Among these methods, head movement-based control has gained increasing attention due to its non-invasiveness, low cost, and ease of use. Head movement signals are typically captured using inertial measurement units (IMUs), gyroscopes, or optical tracking, and are then processed by discrete-state decoders to produce a limited set of commands, such as turning left or right, moving forward, or stopping [1]. However, such individualized biosignal based control approaches and their associated decoders are not able to generate stable outputs (trajectories) particularly when facing the challenge of continuous control of an end-effector based system (e.g. computer cursor) for accurate target acquisition (Fig 1)[11, 12]. This challenge could be more destructive when controlling the endpoint of a physical robotic system where a minimum jerky motion causes failure in tasks such as robotic grasping for a patient with severe motor impairment [14]. Additionally, the required high level of engagement for continuous control of traveled trajectory imposes a high cognitive load on most patients particularly those with brain implants [12]. This cognitive burden was further escalated by the uncertainty of human intention decoders in real-world scenarios of human-robot collaboration for target acquisition, as studied by Dani et al. [15]. Most of the previous research studies on shared human-robot control [16, 17, 18, 19] assumed known or explicitly predicted targets, lacking real-time responsiveness to dynamic user intentions, demands, or changes in the environment’s structure. To this end, we present STREAMS (Self-Training Robotic End-to-end Adaptive Multimodal Shared autonomy), a novel framework using deep reinforcement learning along with a DQN approach, to overcome the barriers of current human-robot shared control approaches in reach-to-grasp objects in 2D workspace by focusing on the following highlighted contributions:
@@ III-A Problem formulation and framework
In our case study, we consider $\tilde{h_{t}}\in\{-1,0,1\}$, which represents decoded values of head movements corresponding to left, neutral, and right, respectively. The action space $\mathcal{A}$ consists of five possible actions: forward, backward, right, left, and hold. We focus on controlling the end-effector position directly, while the positions of other joints in the robotic arm are determined using inverse kinematics. The Q-value function $Q(s_{t},a)$ is approximated by a DQN, which integrates convolutional layers to process image inputs, embedding layers to represent human inputs, and a late fusion mechanism to combine information from both modalities. The output of the fusion mechanism is passed through linear layers to make the final action decision as illustrated in Fig 3.
@@
As mentioned before, biosignal-based control systems, rely on discrete classification outputs from machine learning decoders to translate user intent into robotic actions. However, decoder outputs are inherently prone to misclassification due to signal variability and noise, leading to jerky and unstable control trajectories in real-world applications. Different studies [25, 26, 27, 28] consistently show that for the multiclass motor imagery classification problem using EEG, which is among the most challenging biosignals to classify accurately, typical decoder accuracies range between 70% and 90%, indicating that misclassification rates can be as high as 30% in practical applications. To simulate real-world classification errors in biosignal-based control systems, we introduce a probabilistic flipping function, where ideal control inputs are randomly misclassified with probability $p$ (Equation 5). This approach is inspired by common noise models used in EEG-based brain-computer interfaces and noisy reinforcement learning environments, where errors are modeled as stochastic perturbations to approximate real-world decoder inaccuracies [27, 25]. Unlike continuous noise models, such as Gaussian or uniform noise, the flipping function is more suitable for discrete control signals, such as categorical decisions in EEG, EMG, and IMU-based classification systems [26, 29].
@@
Here, the function $\text{Flip}(h^{ideal}_{t})$ randomly changes $h^{ideal}_{t}$ to one of the other two possible values $\{-1,0,1\}$. In our framework, we simulate three distinct levels of decoder noise by assigning probabilistic misclassification rates of $p=0.2$, $p=0.3$, and $p=0.4$ during the training phase.
The pseudo-code for the training phase of the algorithm is presented in Algorithm 1. The DQN in the STREAMS framework is trained using experience replay, where past experiences $(s_{t},a_{t},r_{t},s_{t+1})$ are stored in a replay buffer, and mini-batches are sampled to update the network parameters $\theta$. A target network with parameters $\theta^{\prime}$ is used to stabilize learning. The loss function involves the difference between the predicted and target Q-values, incorporating the reward $r$ and discount factor $\gamma$. The performance is evaluated on metrics such as task success rate, trajectory efficiency, and robustness to input noise. STREAMS is trained in a simulated environment with a hyperparameter random search using the Adam optimizer and an $\epsilon$-greedy strategy for exploration, with $\epsilon$ decaying from 0.9 to 0.1. Training includes 25,000 episodes and randomization of cup positions to prevent overfitting and enhance generalization across tasks.
@@
Figure 4: Components of the STREAMS framework user study
@@ III-B User study design
This IRB-approved study involved five healthy participants aged 22-36 years. The experimental setup included a Kinova Jaco robotic arm, an RGB camera for workspace perception, an Inertial Measurement Unit (IMU)-based head movement detector, and two randomly selected target objects which are supposed to be grasped by user inputs (see Fig. 4). The head movement detector measured the angle of the head toward the sides and decoded these movements into three discrete control inputs: right, left, and neutral. These user inputs were used to guide the robotic arm to reach two different target cups. The framework was generalized to handle a wide range of target objects including different types of cups, bottles, and containers. The reach-to-grasp operation in a planar 2D space required comparing Manual (pure head movement control) and Assistive (with STREAMS) modes. Each participant completed 40 trials: 20 in Manual mode and 20 in Assistive mode, with 10 trials for each target (left and right) in both modes. Target positions remained constant throughout the experiment to ensure a systematic user study across participants. In this case study, users controlled lateral motion through head movements, while the end-effector constantly moved forward with each step. Grasping was automatically triggered whenever the end-effector was in proximity to the target sufficiently.
Before data collection, participants were trained in both modes and completed practice trials. Mode order was randomized to minimize learning effects. Data collection included video recording, trajectory extraction, targets and end-effector position tracking, and success rate recording. Participants also completed questionnaires on convenience, effort, and satisfaction for each mode. Performance metrics compared trajectory qualities, success rates, and user feedback between Manual and Assistive modes. A paired t-test was used to determine the statistical significance of observed differences.
@@ V Conclusion
This study introduced STREAMS as a task support framework for shared autonomy in assistive robotics, addressing the challenge of generating robust trajectories from unreliable and noisy user inputs for precise robotic grasping. The framework leverages deep reinforcement learning to refine control strategies dynamically, ensuring stable and smooth end-effector movements even under uncertain user input conditions. Our generalizable framework integrated several key components, including an end-to-end self-learning policy that enables real-time interaction with dynamic user intent and adaptation to environmental changes. Unlike many existing approaches that rely on predefined datasets or explicit goal prediction, STREAMS utilizes a synthetic noise model to train its control policies, allowing it to adapt to different biosignal-based input modalities without requiring extensive data collection. This adaptability makes it particularly well-suited for assistive applications where user input variability is high.
STREAMS demonstrated the ability to generate smooth and stable trajectories across various noise levels and user conditions, highlighting its robustness to decoder misclassifications and signal noise. The efficacy of the proposed framework was validated through extensive simulations and real-world user studies, showing significant improvements in task performance, user satisfaction, and control accuracy compared to manual control without task support. The framework consistently outperformed manual methods in both success rate and trajectory efficiency, reducing the cognitive and physical effort required from users to complete complex robotic manipulation tasks. Future research directions could include expanding the framework’scapabilities to handle more complex environments and tasks.